OpenStudy (anonymous):

Find the exact functional value for sin^-1 (sin(5pi/4)) Explain why the answer isn't 5pi/4 So far i have the answer i believe, sin^-1=-sqrt2/2 -pi/2 and pi/2, please check and explain why its not 5pi/4

OpenStudy (anonymous):

it is not \[\frac{5\pi}{4}\] because the range of \[\sin^{-1}\] is \[[-\frac{\pi}{2},\frac{\pi}{2}]\]

OpenStudy (anonymous):

thank you, I was just stuck for that little bit, I appreciate the help.

OpenStudy (anonymous):

\[\sin(-\frac{\pi}{4})=-\frac{\sqrt{2}}{2}\] so that is -the right answer, \[-\frac{\pi}{4}\]